Warm laboratories ready – University of Copenhagen

Forward this page to a friend Resize Print Bookmark and Share

EastGrip - The East Greenland Ice-core Project > Field Diaries Folder > UK diaries 2017 > Warm laboratories ready

30 May 2017

Tuesday 30th of May 2017

Demonstration from the now active physical properties laboratory. Ice crystals in different types of polarized light. The squares are 1 cm on each side.

Theme of the day: Both warm laboratories report ready.

Inside the science trench there are two warm laboratories. In one we will perform continuous stable isotope measurements on sticks of ice core, and in the other we will perform measurements on crystal structure and other physical properties on ice samples and sampling for layers of volcanic ash (tephra) in the ice cores. In total we have five heated cabins in the trenches (one is not manned) and in the picture of today you get a view into four. We also have a freezer cabin, which was featured yesterday.

What we did today:

  1. Mounting lamps and distribution panels in science trench.
  2. Removing holes and bumps in camp snow surface after last days of digging tunnels and caves.
  3. Drilling in two shifts. 24.55 m drilled today until 19.00.
  4. Two shift logging of ice cores. Logging depth 211.89 m.
  5. Working on infrastructure in science trench. One HEMA now running.
  6. In both warm laboratories we are now performing test measurements.
  7. Entrance balloon for entry of new cook’s freezer cave has been moved to extend the entrance and buried again.
  8. German snow sampling.
  9. Japanese aerosol sampling and snow experiments.
  10. Water vapour sampling, methane sampling, boundary layer and meteorology station working.

Ad.3: The drillers have chosen to continue using the 2 m HT drill and RECAP motor section to optimize stability of runs. As we have two RECAP motor sections and two HT drills, this will not affect the coming firn-air drilling campaign. The electronics people continue working on the new drilling electronics, and they have pinpointed a component that is affected by the cold. As soon as they have made modifications, the new electronics will be mounted on the drill.

Weather today: Sunshine. Temp. - 11°C to - 23°C. Wind: 5 kt from SW to SE. Visibility: To horizon.

FL, J.P. Steffensen

Tuesday 30th of May 2017

Views through fogged windows of our warm cabins in the snow caves. From left to right: Drillers Sune and Matthias watch Dennis at the console in the drillers cabin. Loggers Lydia and Emilie warm up in drillers workshop with Trevor. Bruno is mounting the last electricals in the physical properties laboratory. Valerie, Christian and Bruce are activating isotope measurements in the isotope laboratory.

<

← Previous entry   Next entry →